Equivalent capacitance of parallel combination is different from that of series combination.justify


1
Expert's answer
2021-05-24T16:28:28-0400

Assume two capacitor is capacitance

"C_1,C_2"





Series combination

"Q=Q_1+Q_2"

"\frac{1}{C}=\frac{1}{C_1}+\frac{1}{C_2}"

"C=\frac{{C_1}{C_2}}{C_1+C_2}"

When n capacitor connect in series combination "C_1,C_2,C_3........C_n"

Then

"C_{net}=\frac{C}{n}"

Parallel combination

"V=V_1+V_2"

"C=C_1+C_2"

When n capacitor connect in series combination "C_1,C_2,C_3.........C_n"

Then

"C_{net}=nC"